# Point-MoE: Towards Cross-Domain Generalization in 3D Semantic Segmentation via Mixture-of-Experts

### v1

@@ 4 Experiments
We conduct comprehensive experiments to evaluate Point-MoE for cross-domain 3D semantic segmentation. First, we describe baseline methods and experimental setups (Sec. 4.1). Then, we present detailed empirical results under two primary joint training scenarios: (1) indoor-only, and (2) combined indoor-outdoor training (Sec. 4.2). Finally, we provide ablation studies that highlight key design choices of the MoE architecture (Sec. 4.3) and detailed analysis of MoE behaviors (Sec. 4.4).
@@ 4.1 Experimental Setup
Baseline methods. Since our proposed Point-MoE framework is built upon PTv3 [78], we adopt PTv3 and PPT [80] as primary baselines. PPT, originally designed for dataset-specific BatchNorm  [34] and LayerNorm [4] layers and fine-tuning, is adapted here for joint training. Because PPT requires dataset labels at inference time, we follow One-for-All [74] to train a lightweight domain classifier (achieving over 95% accuracy), which we also use for Point-MoE under the domain-aware setting.
To ensure a fair comparison, we begin by benchmarking the dense PTv3 model under the domain-agnostic joint training setting, systematically varying key design factors such as mixed-batch composition and normalization strategy. In this setup, models are trained across multiple domains without access to dataset labels, and no label information is provided at inference time. This setting reflects a realistic deployment scenario where domain metadata is unavailable. For the domain-aware joint training setting, we compare to the PPT + One-for-All baseline. Here, models are trained with access to dataset labels. During inference, we either provide predicted labels using a trained domain classifier or assign a “generic” class to simulate deployment in unknown or mixed domains. This allows us to isolate the value of domain conditioning and assess the robustness of Point-MoE under both domain-aware and domain-agnostic regimes.
Datasets. Experiments are conducted on three indoor datasets—ScanNet [21], S3DIS [2], and Structured3D [88]—and evaluated on both seen and unseen domains, including Matterport3D [12]. We then evaluate the best-performing Point-MoE and baseline configurations under a more challenging indoor-outdoor setting, incorporating outdoor datasets from nuScenes [11], SemanticKITTI [7], and Waymo [65]. Voxel sizes are standardized to 0.02m for indoor scenes and 0.05m for outdoor scenes. Throughout, we denote model capacity using -S (small) and -L (large) suffixes for all architectures, including PTv3, Point-MoE, and PPT.
@@ 4.2 Experimental Results
Table 1: Performance of Joint Training for Indoor 3D Semantic Segmentation. All models are jointly trained on three indoor datasets and evaluated under two settings. In Domain-agnostic joint training, models are evaluated using only the point cloud and the corresponding dataset-specific label set. In Domain-aware joint training, a predicted scene label from a domain classifier is provided at inference time to guide the model. “-G” indicates that model utilizes domain randomization and contains a learnable “generic” domain label during training. All results are reported as mean Intersection-over-Union (mIoU).
Methods Params Activated Mix Layer. ScanNet Structured3D S3DIS Matterport3D
Seen Unseen
Val Val Test Val Val Test
@@
PTv3-ScanNet [78] 46M 46M 75.0 19.8 20.9 8.8 38.9 44.5
PTv3-Structured3D [78] 46M 46M 21.2 81.5 79.4 7.1 21.0 24.0
PTv3-S3DIS [78] 46M 46M 4.9 3.5 5.3 67.6 5.0 3.4
PTv3-Matterport3D [78] 46M 46M 53.1 22.9 23.7 9.8 54.1 56.0
Domain-agnostic joint training
@@
[…]
Domain-aware joint training
[…]
Table 2: Cross-domain 3D Semantic Segmentation across Indoor and Outdoor Scenes. We evaluate the best-performing models from the indoor-only setting in a more challenging joint training setup that spans both indoor and outdoor datasets. This setup better reflects real-world deployment scenarios, requiring models to generalize across diverse domains.
ScanNet Structured3D S3DIS nuScenes SemKITTI Matterport3D Waymo
Indoor Outdoor Indoor Outdoor
Seen Unseen
Methods Params Activated Mix Layer. Val Val Test Val Val Val Val Test Val Test
Domain-agnostic joint training
[…]
Domain-aware joint training
[…]
@@
Single-dataset training. We begin by evaluating the standard setting in which models are trained and evaluated on individual datasets. Note we train end evaluate PTv3 with language-guided classification. As shown in Tab. 1, PTv3 achieves strong in-domain performance, but its performance drops significantly on unseen domains. For instance, a PTv3 model trained on S3DIS achieves only 3.5% on Structured3D and 4.9% on ScanNet. While domain-specific models are not expected to generalize across domains or categories, these results underscore the brittleness of single-dataset training. This motivates the need for cross-domain training strategies that expose the model to diverse data and promote broader generalization.
Domain-agnostic joint training. In the indoor-only setting, Point-MoE outperforms all variants of PTv3 in most domains and performs competitively on S3DIS. Referring to Tab. 1, Point-MoE-S achieves 75.4% on ScanNet and 69.7% on S3DIS, outperforming PTv3-S by +4.1% and +13.8%, respectively. Remarkably, Point-MoE-S activates only 52M parameters compared to 97M in PTv3-L, yet still surpasses PTv3-L by +0.8% on ScanNet and +8.4% on S3DIS. In addition, Point-MoE-L reaches 75.8% on ScanNet and 73.0% on Structured3D, significantly ahead of PTv3-L. On the unseen Matterport3D domain, Point-MoE-L obtains 41.6% on validation set and 47.8% on test set, outperforming PTv3-L by +1.4% and +1.2% respectively while activating only 59M parameters at inference. In the more challenging indoor and outdoor setting, Point-MoE-L again leads across the board with a few being competitve. Referring Tab. 2, Point-MoE-L achieves 69.2% on S3DIS validation and 70.2% on test split, outperforming the best dense baseline by +5.4% and +6.3%, respectively. On outdoor datasets, Point-MoE-L achieves 72.9% on nuScenes exceeding the dense baseline by +2.6%. Notably, on the unseen datasets, Point-MoE-L outperforming all dense PTv3 variants by a clear margin. These gains are achieved while activating nearly half the parameters of PTv3-L, underscoring the efficiency and scalability of sparse expert routing.
Domain-aware joint training. We further evaluate Point-MoE in the domain-aware setting, where models are trained with access to dataset labels. In the indoor-only setting, both domain-aware Point-MoE and its generic variant (Point-MoE-G) surpass all strong baselines: variants of PPT [80] + One-for-All [74]. As shown in Tab. 1, Point-MoE-G-L achieves 76.7% on ScanNet, 68.0% on Structured3D, and 67.0% on S3DIS, outperforming the best PPT + One-for-All baseline by +2.0%, +3.1%, and +2.3%, respectively. On the unseen Matterport3D dataset, it reaches 41.8% on the validation split and 47.8% on the test set, again outperforming all baselines. Notably, using only 20% of training data tagged as “generic” does not degrade performance on seen domains.
In the indoor-and-outdoor setting, Point-MoE-G-L continues to outperform all baselines across the board. Referring to Tab. 2, it achieves 77.0% on ScanNet, 72.1% and 74.2% on Structured3D (val/test), 62.4% on S3DIS and 72.7% on nuScenes—surpassing the best baseline by +0.3%, +4.6%, +4.1%, +1.0% and +0.5%, respectively. On the unseen domains, including Matterport3D and Waymo, Point-MoE-G-L delivers the strongest generalization. Crucially, it performs on par with domain-aware Point-MoE on seen datasets while achieving superior accuracy on unseen ones. These results validate that domain randomization with a generic label can enhance robustness without compromising in-domain performance. We still observe a performance gap between cross-domain and single-dataset training in cases like Structured3D [88]. Single-dataset models can adapt to domain-specific cues and distributions, while Point-MoE must balance capacity across domains. As a result, its performance may trail on large, homogeneous datasets, but remains competitive overall due to its generalization ability.
4.3 MoE Design Ablation
@@
Table 3: Ablation of Point-MoE Design and Hyperparameters. We conduct a systematic ablation on key design and hyperparameter choices of Point-MoE across four indoor multi-domain datasets: ScanNet (Scan), Structured3D (S3D), S3DIS, and Matterport3D (Mat).
[…]
@@
(b) Top-$k$ expert selection. Best performance is achieved with $k{=}2$ experts.
[…]
@@
[…]
dec. 74.0 64.9 66.6 38.1
[…]
(d) MoE position. Inserting MoE in all layers yields the best results.
[…]
@@
(h) Number of experts. More experts improve performance.
[…]
4 74.0 64.9 66.6 38.1
[…]
(i) Batch size. Larger batchsize improves performance.
We investigate the impact of MoE design choices and hyper-parameters choices in indoor cross-domain 3D semantic segmentation. All experiments are conducted with ScanNet [21], Structured3D [88] and S3DIS [2] and evaluate without providing dataset label on seen dataset and also evaluate on unseen dataset: Matterport3D [12]. Results are presented in Tab. 3.
Auxiliary load-balancing loss. The auxiliary load-balancing loss is designed to encourage uniform expert usage and prevent expert collapse [54, 76, 43]. We ablate its effect by training models with varying strengths of this loss in Tab. 3. Interestingly, we find that removing the auxiliary loss entirely leads to consistently better performance across domains. In contrast, stronger auxiliary losses result in significant degradation. We hypothesize that this is due to the inherent imbalance in the distribution of samples across domains in 3D point cloud datasets—a phenomenon also observed by ChartMoE [82].
Effect of activated experts. Tab. 3 compares the performance of configurations with 1, 2 and 3 activated experts, finding diminishing return when increasing number of activated experts. We find that activating two experts yields the best performance across domains. This analysis suggests that merely increasing the number of activated experts does not guarantee improved performance.
MoE position. We ablate the placement of MoE layers to understand where expert specialization is most effective. Empirically, we observe greater expert diversity when MoE is applied in the decoder compared to the encoder. In Tab. 3, we compare two configurations: applying MoE layers only in the decoder, versus applying them in both the encoder and decoder. Results show that applying MoE to both stages consistently yields better performance, suggesting that expert specialization benefits from exposure to both encoder and decoder.
@@
Effect of scaling model with MoE. Tab. 3 compares Point-MoE with 4 and 8 experts with activated experts fixed at 2. Increasing the number of experts from 4 to 8 leads to improved performance on seen domains, and yields a modest gain of 0.4% on the unseen domain. Tab. 3 ablates the expert width by varying the intermediate size of the MLP (e.g., 1H, 2H). The results show a mixed trend: wider experts improve performance on ScanNet and Matterport3D, but degrade results on Structured3D and S3DIS. Finally, we investigate the effect of adding shared experts [63, 72, 22], which are applied universally without gating. We find that disabling shared experts consistently leads to better accuracy and domain specialization across all benchmarks, as reported in Tab. 3.
@@
Ablation of hyper-parameters choices. We provide justification for our selected hyperparameters. In Tab. 3, BatchNorm [34] consistently outperforms other normalization methods [4, 85] in the cross-domain setting. Tab. 3 shows that ReLU [1] slightly outperforms SiLU [27]. Additionally, Tab. 3 demonstrates that importance of batch size in MoE training. Specifically, the batch size must exceed the number of experts to avoid performance degradation. In general, larger batch sizes lead to improved performance across domains.

### v2

@@ 4 Experiments
We conduct comprehensive experiments to evaluate Point-MoE for 3D semantic segmentation. First, we describe baseline methods and experimental setups (Sec. 4.1). Then, we present detailed empirical results under two primary joint training scenarios: (1) indoor-only, and (2) combined indoor-outdoor training (Sec. 4.2). Finally, we provide ablation studies that highlight key design choices of the MoE architecture (Sec. 4.3) and detailed analysis of MoE behaviors (Sec. 4.4).
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on three indoor benchmarks: ScanNet [19], S3DIS [1], and Structured3D [95], and additionally evaluate zero-shot on Matterport3D [10]. Specifically, we train a single model on ScanNet, S3DIS, and Structured3D, and directly evaluate its performance on both seen and unseen datasets via the language-guided head using the same checkpoint. We further test the best-performing Point-MoE architectures and baseline configurations in a more challenging indoor and outdoor joint setting that incorporates nuScenes [9] and SemanticKITTI [5] as training data, and Waymo [67] for zero-shot evaluation. The same evaluation protocol is used across both settings.
Baselines. Since Point-MoE is built on PTv3 [81], we adopt PTv3 and PPT [82] as our primary baselines. PPT, originally designed with dataset-specific normalization and per-dataset fine-tuning, is adapted here to the joint training setting. Because PPT requires dataset labels at inference, we follow One-for-All [75] and train a lightweight dataset classifier (achieving $\geq$95% accuracy) to predict the source dataset for an input sample at test time (Sec. B.2).
Evaluation and Implementation Details. To ensure fair comparison, we re-benchmark PTv3 under our unified protocol by varying batch composition and normalization, which yields an average improvement of 22 mIoU (Sec. B.3). We also compare against PPT combined with One-for-All, which uses ground-truth dataset labels during training and predicted dataset labels during inference. Model capacity is denoted with -S (small) and -L (large) suffixes for all architectures, including PTv3, Point-MoE, and PPT. PTv3-S and PPT-S use 14 encoder and 8 decoder layers, while their -L variants use 18 encoder and 12 decoder layers; in contrast, both Point-MoE-S and -L have 14 encoder and 8 decoder layers, with -S using 4 experts per MoE layer and -L using 8. The expert hidden dimension is set to 1× the input dimension in Point-MoE-S and 2× in Point-MoE-L. For details, please refer to Tab. 10 in the supplementary material.
@@ 4.2 Experimental Results
Table 1: Results on seen datasets. Results are reported in mIoU. Point-MoE achieves the best average performance without using domain labels. Methods marked with use dataset labels during training. Models trained on a single dataset but evaluated on unseen datasets are in gray. (a) Indoor-only joint training on ScanNet [19], Structured3D [95], and S3DIS [1]. (b) Joint training across indoor and outdoor datasets: nuScenes [9] and SemanticKITTI [5].
[…]
(a) Indoor-only joint training
Methods Params Activated ScanNet Structured3D S3DIS Avg.
Val Val Test Area5
@@
PTv3-ScanNet [81] 46M 46M 75.0 19.8 20.9 8.8 31.1
PTv3-Structured3D [81] 46M 46M 21.2 81.5 79.4 7.1 47.3
PTv3-S3DIS [81] 46M 46M 4.9 3.5 5.3 67.6 20.3
Multi-dataset Joint Training
OA-CNNs [52] 52M 52M 71.6 59.0 58.6 60.6 62.5
PTv3-S [81] 46M 46M 71.3 56.6 55.9 69.4 63.3
PTv3-L [81] 97M 97M 72.2 56.0 58.1 67.4 63.4
PPT-S [82, 75] 47M 47M 74.7 64.2 65.5 64.7 67.3
PPT-L [82, 75] 98M 98M 74.8 64.7 65.9 65.0 67.6
Point-MoE-S 59M 52M 74.5 69.0 69.9 68.0 70.4
Point-MoE-L 100M 59M 76.0 69.5 70.8 69.5 71.5
@@
(b) Indoor and Outdoor joint training
Methods Params Activated ScanNet Structured3D S3DIS nuScenes SemKITTI Avg.
Val Val Test Area5 Val Val
PTv3-L [81] 97M 97M 74.6 64.1 63.9 68.0 69.6 62.9 67.2
PPT-L [75, 82] 98M 98M 76.7 67.5 66.0 61.9 70.8 66.9 68.3
Point-MoE-L 100M 59M 77.4 69.5 70.2 68.1 72.9 66.4 70.8
Table 2: Zero-shot results on unseen datasets. Point-MoE consistently achieves the strongest generalization across unseen datasets in the zero-shot setting. Methods marked with use domain labels during training.
@@
(a) Indoor-only joint training
Method Matterport3D Average
Val Test
PTv3-L 39.0 45.4 42.2
PPT-L 27.8 26.3 27.1
Point-MoE-L 41.0 46.7 43.9
@@
(b) Indoor and Outdoor joint training
Method Matterport3D Waymo Average
Val Test Val Test
PTv3-L 39.1 45.9 21.9 23.2 32.5
PPT-L 24.9 23.1 16.3 16.7 20.3
Point-MoE-L 41.3 49.6 23.7 25.3 35.0
@@
Single-dataset training generalizes poorly. We begin by evaluating the standard setting where models are trained and evaluated on individual datasets. Note that in our setup PTv3 is trained and evaluated with CLIP-guided classification in an open-set setting, consistent with the multi-dataset regime [82], whereas the original PTv3 paper reports results under standard closed-set softmax label classification [81]. This difference makes our scores not directly comparable to the numbers in the original PTv3 paper. As shown in Tab. 11, PTv3 achieves strong results on seen datasets, but its performance drops significantly on zero-shot datasets. For instance, a PTv3 model trained on S3DIS achieves only 3.5% on Structured3D and 4.9% on ScanNet. While dataset-specific models are not expected to generalize across datasets or categories, these results underscore the brittleness of single-dataset training. This motivates the need for multi-dataset joint training strategies that expose the model to diverse data and promote broader generalization.
Point-MoE surpasses all baselines on seen datasets, without dataset labels during training. Point-MoE dispenses with dataset identifiers during training yet yields consistent gains across indoor benchmarks (Tab. 11); both Point-MoE-S and Point-MoE-L variants improve on ScanNet, Structured3D, and S3DIS, with Point-MoE-L attaining a new state-of-the-art average of $71.5$ mIoU. Despite joint training over multiple datasets, Point-MoE matches or exceeds single-dataset specialists, indicating effective expert routing rather than overfitting to any one source, or significantly exhibiting the negative transfer effects noted in PPT [82]. In the more challenging indoor and outdoor joint training setting (Tab. 11), Point-MoE maintains its advantage across datasets and, on average, surpasses PTv3-L by $3.55$ mIoU and PPT-L by $2.45$ mIoU, demonstrating that label-free expert routing scales to heterogeneous datasets while simplifying the training pipeline. Moreover, performance on the seen indoor datasets remains closely aligned between indoor-only training and indoor-outdoor joint training, suggesting that Point-MoE preserves seen datasets accuracy even when subsequently trained jointly with outdoor data. This generalization is achieved implicitly in the model, without requiring dataset labels for incoming samples or maintaining explicit per-dataset normalization parameters. Note, since we use a lightweight dataset classifier, PPT’s performance differs from what is reported in the original paper, where oracle dataset identifiers are provided at inference. Finally, because only a subset of experts is activated, Point-MoE achieves a favorable accuracy–compute trade-off, cutting compute by $30.9\%$ (265.7 vs. 384.4 GFLOPs) and peak VRAM by $19.0\%$ (33.3 vs. 41.1 GiB) relative to PPT-L (Tab. 5). We analyze the emergent dataset-awareness and routing behavior in Sec. 4.
Point-MoE outperforms all baselines consistently across zero-shot datasets. As shown in Tab. 22 and Tab. 22, Point-MoE-L achieves the strongest zero shot performance, including on Matterport3D under indoor only training and on both Matterport3D and Waymo under indoor and outdoor training, without using dataset labels during training. We believe a key reason is that label free expert routing encourages specialization based on the underlying semantics and geometry of the input rather than on dataset identity, resulting in more stable behavior under distribution shift. By contrast, PPT-L relies on explicit dataset labels during training. While this provides useful supervision on seen datasets, it can also encourage the model to depend too heavily on dataset specific cues. As a result, when a dataset is held out and its label is unavailable or no longer aligned with the training distribution at test time, this dependence becomes brittle and can hurt generalization. Additional analysis of Point-MoE’s zero shot adaptation is provided in Sec. 4.4.
@@
Table 3: Ablations of Point-MoE designs and hyperparameters. We conduct a systematic ablation on key design and hyperparameter choices of Point-MoE across four indoor multi-domain datasets: ScanNet (Scan), Structured3D (S3D), S3DIS, and Matterport3D (Mat).
$\alpha$ Scan S3D S3DIS Mat
0 74.5 66.9 68.6 40.1
@@
(b) Top-$k$ expert selection. Best performance is achieved with k=2k{=}2 experts across datasets.
norm Scan S3D S3DIS Mat
BN. 74.5 66.9 68.6 40.1
@@
case Scan S3D S3DIS Mat
FFN 72.6 66.1 66.9 39.2
Proj. 74.5 66.9 68.6 40.1
(d) MoE position. Replacing the projection layer performs better.
act. Scan S3D S3DIS Mat
SiLU 73.6 65.7 64.9 40.1
ReLU 74.5 66.9 68.6 40.1
@@
num. Scan S3D S3DIS Mat
4 74.5 66.9 68.5 40.1
8 73.6 68.7 68.5 40.4
(h) Number of experts. More experts improve performance.
batch Scan S3D S3DIS Mat
4 74.0 64.9 66.6 38.1
6 74.3 66.9 68.6 40.1
(i) Batch size. Larger batchsize improves performance.
4.3 MoE Design Ablation
We investigate the impact of MoE design choices and hyper-parameters choices in indoor multi-dataset 3D semantic segmentation. Results are presented in Tab. 3. Note we are not using precise evaluator so results are not directly comparable to Tab. 1. Please refer to details in Sec. B.3 in the supplementary material.
Effect of load-balancing loss. The auxiliary load-balancing loss is designed to encourage uniform expert usage and prevent expert collapse [56, 78, 42]. We ablate its effect by training models with varying strengths of this loss in Tab. 3. Interestingly, we find that removing the auxiliary loss entirely leads to consistently better performance across datasets. In contrast, stronger auxiliary losses result in significant degradation. We hypothesize that this is due to the inherent imbalance in the distribution of samples across datasets in 3D point cloud datasets—a phenomenon also observed by ChartMoE [87].
Effect of activated experts. Tab. 3 compares the performance of configurations with 1, 2 and 3 activated experts, finding diminishing return when increasing number of activated experts. We find that activating two experts yields the best performance across datasets. This analysis suggests that merely increasing the number of activated experts does not guarantee improved performance.
Table 4: Ablations for mixed-dataset training. “Mix” indicates that each training batch combines samples from multiple datasets.
Methods Mix Scan S3D S3DIS Mat Average
PTv3 48.4 57.8 44.1 27.8 44.5
@@ 4.3 MoE Design Ablation
Effect of MoE position. We compare placing MoE in the feed-forward network (FFN) versus in the attention output projection $W_{o}$, holding all other components fixed. Projection-MoE consistently outperforms FFN-MoE in Tab. 3. We hypothesize that $W_{o}$ is the fusion point where multi-head outputs are recombined before the next normalization, so experts placed there can route on richer, head-aggregated signals that preserve cross-dataset geometry cues, while leaving $W_{q},W_{k},W_{v}$ shared and stable; in contrast, FFN experts operate on post-normalized activations where such dataset or geometry cues are more attenuated. This empirical gain leads us to adopt projection-MoE as the default. Similar observations also appear in recent work: e.g. UMoE [89] finds that using MoE for the output (and value) projections yields strong compute savings with high performance.
@@
Effect of model scaling with MoE. Tab. 3 compares Point-MoE with 4 and 8 experts with activated experts fixed at 2. Increasing the number of experts from 4 to 8 leads to improved performance on seen and zero-shot datasets. Tab. 3 ablates the expert width by varying the intermediate feature dimension and the results show a mixed trend: wider experts improve performance on ScanNet and Matterport3D, but degrade results on Structured3D and S3DIS. Finally, we investigate the effect of adding shared experts [65, 73, 20]. We find that disabling shared experts consistently leads to better accuracy and dataset specialization across all benchmarks, as reported in Tab. 3.
@@
Effect of hyper-parameter choices. We provide justification for our selected hyper-parameters. In Tab. 3, BatchNorm [29] consistently outperforms other normalization methods [3, 91] in the multi-datasets setting. Tab. 3 shows that ReLU [50] slightly outperforms SiLU [23]. Additionally, Tab. 3 demonstrates that importance of batch size in MoE training. In general, larger batch sizes lead to improved performance across datasets.
Mixed dataset training. In joint multi-dataset training, forming each minibatch with samples drawn from multiple datasets is consistently beneficial. Although there is no explicit batch-level interaction module, samples from different datasets still influence one another through shared normalization statistics and, for Point-MoE, through competition over a common expert pool. As shown in Tab. 4, this mixed-dataset training strategy improves the average mIoU of all backbones: PTv3 +8.3+8.3, PPT +2.9+2.9, and Point-MoE +16.6+16.6 mIoU on average. For dense backbones such as PTv3 and PPT, mixing exposes the model to more diverse domain statistics within each optimization step, which appears to regularize training and reduce overfitting to dataset-specific feature distributions. To make this comparison fair for PPT, we re-implemented its normalization module so that mixed-dataset batching is supported. The effect is most pronounced for Point-MoE, where diverse minibatches promote expert specialization across datasets, stabilizing routing and improving generalization.
Figure 2: t-SNE visualization of feature clustering. The first three columns show that the decoders of both PPT and PointMoE have better separation between datasets in their decoder representations, than vanilla PTv3. The rightmost column highlights how Point-MoE generalizes to zero-shot datasets, with Matterport3D features aligning closely with ScanNet, which is indeed semantically similar.